# Machine Learning and the Physical Sciences

### v1

@@ Abstract
Reduced-order models (ROMs) can efficiently simulate high-dimensional physical systems, but lack robust uncertainty quantification methods. Existing approaches are frequently architecture- or training-specific, which limits flexibility and generalization. We introduce a post hoc, model-agnostic framework for predictive uncertainty quantification in latent space ROMs that requires no modification to the underlying architecture or training procedure. Using conformal prediction, our approach estimates statistical prediction intervals for multiple components of the ROM pipeline: latent dynamics, reconstruction, and end-to-end predictions. We demonstrate the method on a latent space dynamical model for cloud microphysics, where it accurately predicts the evolution of droplet-size distributions and quantifies uncertainty across the ROM pipeline.
@@ 1 Introduction
Latent space reduced-order modeling learns a compact representation of high-dimensional physical dynamics in a lower-dimensional latent space. These models are valuable for scientific applications where the governing physics are partially known or computationally prohibitive to resolve. For example, accurately resolving clouds and precipitation in an atmospheric simulation would require tracking high-dimensional droplet-size-distributions (DSDs), a longstanding parametric challenge in climate and weather modeling known as “cloud microphysics.” Error and uncertainty in microphysics parameterizations is typically not quantified, but is believed to be a dominant source of uncertainty in future climate projections [morrison_confronting_2020].
Latent space ROMs have proven effective in efficiently simulating related complex fluid mechanical systems (e.g. [choi_sns_2020, fries_lasdi_2022], yet convincing practitioners of their reliability is challenging due to the lack of unified and robust uncertainty quantification (UQ) frameworks. Existing UQ methods for latent space dynamical models are often tied to specific architectures [Bonnet2023, cheng2025latentspaceenergybasedneural, YONG2025117638], require expensive training [Simpson2024-ie], or make parametric assumptions [iakovlev2023latentneuralodessparse, Kingma_2019].
We present a model-agnostic, post hoc framework for predictive UQ in latent-space ROMs that quantifies uncertainty on reconstruction, latent dynamics, and end-to-end predictions without altering the base architecture or training procedure. Our approach utilizes conformal prediction (CP), a distribution-free method that produces statistically valid prediction intervals – a first for latent space ROMs. We demonstrate this UQ pipeline on a cloud microphysics ROM trained to predict the evolution of cloud DSDs during coalescence and the formation of precipitation, producing reliable UQ estimates that allow practitioners to rigorously evaluate individual components of the ROM architecture.
@@ 2 Proposed UQ framework
A latent space dynamical ROM consists of a data space, $\mathcal{X}\subseteq\mathbb{R}^{d}$, and a latent space, $\mathcal{Z}\subseteq\mathbb{R}^{D}$, where $D\ll d$; an encoder $E:\mathcal{X}\rightarrow\mathcal{Z}$ and a decoder $D:\mathcal{Z}\rightarrow\mathcal{X}$; and a dynamical system model $F:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{Z}$ defined on the latent space, where $\mathcal{T}\subseteq\left[0,T\right]$ for some final time $T>0$. (See Figure 1.)
Figure 1: A generic latent space dynamical model with fixed dynamical time-step $\Delta t$. Figures 1a and 1b show the reconstruction and dynamics sub-models, respectively, that comprise the end-to-end model architecture shown in Figure 1c.
We consider the setting where we observe $n$ independent time-evolving realizations of a physical system in $\mathcal{X}$: $\{x_{t}^{(j)}\}_{t\in\mathcal{T}}$ for $j=1,\dots,n$. The proposed pipeline applies regardless of whether model components are trained separately or end-to-end. Hence, at each fixed time $t\in\mathcal{T}$ and across samples $j=1,\dots,n$, we compute predictive uncertainties in the components of a latent space dynamical model: 1) reconstruction from the latent space, 2) dynamics in the latent space, and 3) the entire pipeline combined (end-to-end), as shown in Figure 1.
@@
Rather than assuming errors are symmetric around the predictions, we follow a two-sided approach to construct prediction intervals using the $\alpha/2$ and $1-\alpha/2$ quantiles of the signed residuals [barber_predictive_2021]. This modification produces asymmetric prediction bands with valid $1-\alpha$ coverage—a crucial feature for DSDs, as these are non-negative and often contain values near zero, allowing prediction intervals to reflect physical uncertainty better. See Appendix B for further details.
@@ 3 Application to Cloud Microphysics ROM
We demonstrate the UQ pipeline with a cloud microphysics ROM trained on DSDs from particle-based simulations of warm-rain coalescence. Coalescence consists of smaller cloud droplets combining to form larger ones, eventually leading to precipitation [khain_representation_2015]. As shown in Figure 2, droplet mass shifts from smaller to larger size bins under coalescence, and the strong nonlinearity of the process drives the emergence and disappearance of multiple modes.
@@
New parameterizations of cloud microphysics must not only outperform traditional schemes in accuracy and efficiency, but also quantify structural errors and parametric uncertainties that currently hinder the accuracy of large-scale models. Nonlinear ROMs offer a more flexible alternative to bulk parameterizations, which impose restrictive modeling assumptions [kessler_distribution_1969, khairoutdinov_new_2000, seifert_double-moment_2001], and to linear latent space ROMs, which could require inefficiently high-dimensional representations to capture DSD coalescence dynamics accurately [lin_enhanced_2025, romor_non-linear_2023, wan_evolve_2023]. Hence, to model droplet coalescence, we combine an autoencoder (AE) for nonlinear dimensionality reduction with parsimonious latent space ODEs, based on the Sparse Identification of Nonlinear Dynamics (SINDy) technique [brunton_discovering_2016, champion_data-driven_2019]. (For details of the AE-SINDy architecture and training, see Appendix C and the associated code repository. Here, we focus on the UQ pipeline and results.)
Our UQ pipeline is uniquely suited to address both structural uncertainty in the autoencoder compression of DSDs to the latent space and parametric uncertainty in the identified ODEs, tracing how these uncertainties propagate into DSD coalescence predictions. The examples in Figure 2 demonstrate uncertainty for one sample DSD at one time, while Figure 3 aggregates prediction intervals across samples to assess uncertainty arising from different parts of the model as functions of time. Our analysis highlights the ability of the presented UQ pipeline to characterize data-driven ROMs by highlighting the specific scales and processes where model improvements are most necessary.
@@
Figure 2: Initial and final states for three sample DSD trajectories from the dataset, evolved using the AE-SINDy pipeline with empirical prediction intervals provided via CV+ conformal predictions (using $k=20$ folds) at varying nominal coverage levels.
@@
Figure 3: Weighted average prediction interval width vs. time, computed by applying vanilla CP, split CP (using a 6060-2020-2020 train-validation-test split), and CV+ conformal predictions (using $k=20$ folds), respectively, to the indicated components of the AE-SINDy pipeline at varying nominal coverage levels $1-\alpha$. We assess this mean predictive uncertainty by averaging the widths of prediction intervals across samples as a function of time for various miscoverage levels $\alpha$, using integrated totals for DSD predictions and averages for latent variables.
@@ 4 Results & Discussion
Figure 3 illustrates not only predictive accuracy on unseen data but also qualitative patterns of uncertainty propagation across the AE-SINDy architecture. As anticipated, reconstruction uncertainty intervals, which characterize the autoencoder alone, are consistent across time. The latent dynamics exhibit rapid incipient growth in uncertainty that slows over time, reflecting the challenging cloud-to-rain transition before droplets settle into a rain-dominant coalesced state. By contrast, the full ROM produces nearly linear growth in the end-to-end predictive errors. This highlights a key advantage of component-wise uncertainty analysis in ROMs: we detected how latent errors are effectively “smoothed out” by the autoencoder, yielding linear error growth in the final predictions.
@@
While all conformal prediction methods achieved near-nominal coverage (see Appendix A), CV+ yielded wider average intervals to produce more reliable coverage, but at a higher computational cost. CV+ requires retraining the surrogate model $k$ times, compared to just once for vanilla or split conformal. The cost of training the surrogate model will therefore determine practical choices for CP techniques in future applications.
@@ 5 Limitations & Future Directions
The agreement across CP methods and variability in empirical coverages was notably better for the reconstruction and end-to-end network predictions than for latent dynamical predictions (cf. Figure 3). This is likely due to greater variance in prediction fidelity across times and variables—as well as a wider range of magnitudes overall—in latent predictions compared to normalized DSD predictions. Increasing training data—for instance, by widening the simulation domain used to generate the dataset (see Appendix C.1) or increasing its resolution—could reduce this variability, yielding more consistent CP intervals and improving coverage accuracy on the test set.
@@
This work was performed in part under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory (LLNL) under Contract DE-AC52-07NA27344 and supported by the Laboratory Directed Research and Development Program (LDRD), project number 25-ERD-045. J. Katona was partially supported by a stipend and teaching fellowship from the Yale Graduate School of Arts and Sciences. The authors have declared that none of them has any competing interests. Released under IM number LLNL-CONF-2010541.
The authors would like to thank LLNL—particularly Livermore Computing—for their support and for providing high-performance computing resources that enabled the simulations, data analyses, and figure generation in this study.
@@ Appendix A Validation of empirical prediction intervals
Tables 1 and 2 give summary statistics—means and standard deviations, and medians, respectively—for the nominal coverages in the outputs for different subsets of the network: autoencoder/reconstruction alone, latent dynamical model alone, and the entire end-to-end network predictions. These statistics are averaged over all output variables—64 bins for DSD data or 4 latent variables—and times—61 timesteps at 10 second intervals.
@@
Table 1: Empirical coverage (in %, given as empirical mean $\pm$ empirical standard deviation across all times and output coordinates) for prediction intervals at marginal coverage levels 90%, 95%, 98%, and 99%, computed using the CP methods indicated. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Empirical medians can be found in Table 2.
CP Method Sub-model $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Vanilla Reconstruction 88.56 $\pm$ 3.15 93.85 $\pm$ 2.30 96.98 $\pm$ 1.63 98.01 $\pm$ 1.35
Vanilla Latent dynamics 86.00 $\pm$ 12.70 91.80 $\pm$ 12.66 95.39 $\pm$ 12.57 96.88 $\pm$ 12.64
Vanilla End-to-end 88.62 $\pm$ 3.35 93.71 $\pm$ 2.58 96.79 $\pm$ 1.97 97.93 $\pm$ 1.60
Split Reconstruction 87.70 $\pm$ 4.31 92.87 $\pm$ 3.61 96.10 $\pm$ 2.56 97.34 $\pm$ 2.17
Split Latent dynamics 86.93 $\pm$ 11.80 92.05 $\pm$ 12.22 95.14 $\pm$ 12.44 96.48 $\pm$ 12.54
Split End-to-end 87.38 $\pm$ 4.58 92.78 $\pm$ 3.56 96.26 $\pm$ 2.53 97.47 $\pm$ 1.94
CV+ Reconstruction 89.03 $\pm$ 3.09 94.36 $\pm$ 2.23 97.35 $\pm$ 1.61 98.28 $\pm$ 1.36
CV+ Latent dynamics 91.08 $\pm$ 12.96 95.67 $\pm$ 12.73 97.55 $\pm$ 12.71 98.01 $\pm$ 12.68
CV+ End-to-end 90.47 $\pm$ 3.59 95.22 $\pm$ 2.43 97.66 $\pm$ 1.70 98.49 $\pm$ 1.32
@@
Table 2: Empirical coverage (in %, given as the empirical median across all times and output coordinates) for prediction intervals at marginal coverage levels $1-\alpha=$90%, 95%, 98%, and 99%, computed using the CP methods indicated. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Empirical means and empirical standard deviations can be found in Table 1.
CP Method Sub-model $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Vanilla Reconstruction 88.71 93.55 96.77 98.39
Vanilla Latent dynamics 89.52 94.35 96.77 99.19
Vanilla End-to-end 88.71 93.55 97.58 98.39
Split Reconstruction 87.10 93.55 95.97 97.58
Split Latent dynamics 88.71 94.35 96.77 98.39
Split End-to-end 87.10 92.74 96.77 97.58
CV+ Reconstruction 89.11 94.35 97.58 98.39
CV+ Latent dynamics 93.55 97.58 100.00 100.00
CV+ End-to-end 91.13 95.16 97.58 99.19
@@
Across all CP methods, the empirical coverage was generally close to nominal levels, indicating good calibration. While split conformal is theoretically more accurate than vanilla conformal [angelopoulos2021gentle, lei2018distribution], for both reconstruction and end-to-end outputs, vanilla and split conformal performed similarly, with mean coverages typically within 1%–2% of the target and relatively low variability and standard deviations under $\sim$4%. By contrast, CV+ consistently achieved slightly higher accuracy, producing empirical coverages that were closer to the nominal rates in both mean and median, especially at higher confidence levels. This improvement was most apparent in the end-to-end model, where CV+ coverage levels tracked the nominal ones more tightly than vanilla or split conformal did.
The latent dynamical model posed the greatest challenge for all CP methods, with notably higher variability in empirical coverage. Standard deviations exceeded 12% for vanilla and split conformal across all levels, reflecting unstable coverage across output variables and times. That being said, CV+ was partially able to mitigate this issue, yielding higher mean and median coverages that sometimes matched or exceeded the target levels. For instance, CV+ displayed a median coverage of 93.55% vs. 90% at the 90% nominal level, and even full coverage at the 98% and 99% levels. These results suggest that, while CP methods reliably capture uncertainty for reconstructions and end-to-end outputs, consistently quantifying predictive uncertainty for latent dynamics remains challenging, though CV+ provides the most robust coverage.
@@
To test more specifically for consistency in the predictive intervals across different CP methods, we also refer the reader to Figure 3. For all three subsets of the network, the predictive errors became less consistent as $\alpha\rightarrow 0$. This is because the convergence of empirical quantiles to their true values depends strongly on the miscoverage rate $\alpha$ [serfling1980chapter2, vaart1998chapter21]. In particular, for CP, predictive intervals at smaller $\alpha$ require larger calibration sets to stabilize because extreme quantiles converge more slowly, whereas more central quantiles yield more stable intervals with fewer samples [lei2018distribution, vovk_algorithmic_2022].
Appendix B Tailwise conformal prediction intervals

### v2

# Uncertainty Quantification for Reduced-Order Surrogate Models Applied to Cloud Microphysics

@@ Abstract
Reduced-order models (ROMs) can efficiently simulate high-dimensional physical systems but lack robust uncertainty quantification methods. Existing approaches are frequently architecture- or training-specific, which limits flexibility and generalization. We introduce a post hoc, model-agnostic framework for predictive uncertainty quantification in latent-space ROMs that requires no modification to the underlying architecture or training procedure. Using conformal prediction, our approach estimates statistical prediction intervals for multiple components of the ROM pipeline: latent dynamics, reconstruction, and end-to-end predictions. We demonstrate the method on a latent-space dynamical model for cloud microphysics, where it accurately predicts the evolution of droplet-size distributions and quantifies uncertainty across the ROM pipeline.
@@ 1 Introduction
Latent-space reduced-order modeling learns a compact representation of high-dimensional physical dynamics in a lower-dimensional latent space. These models are valuable for scientific applications where the governing physics are partially known or computationally prohibitive to resolve. For example, accurately resolving clouds and precipitation in an atmospheric simulation would require tracking high-dimensional droplet-size distributions (DSDs), a longstanding parametric challenge in climate and weather modeling known as “cloud microphysics.” Error and uncertainty in microphysics parameterizations is typically not quantified, but is believed to be a dominant source of uncertainty in future climate projections [morrison_confronting_2020].
Latent-space ROMs have proven effective in efficiently simulating related complex fluid mechanical systems (e.g. [choi_sns_2020, fries_lasdi_2022]), yet convincing practitioners of their reliability is challenging due to the lack of unified and robust uncertainty quantification (UQ) frameworks. Existing UQ methods for latent-space dynamical models are often tied to specific architectures [Bonnet2023, cheng2025latentspaceenergybasedneural, YONG2025117638], require expensive training [Simpson2024-ie], or make parametric assumptions [iakovlev2023latentneuralodessparse, Kingma_2019].
We present a model-agnostic, post hoc framework for predictive UQ in latent-space ROMs that quantifies uncertainty on reconstruction, latent dynamics, and end-to-end predictions without altering the base architecture or training procedure. Our approach utilizes conformal prediction (CP), a distribution-free method that produces statistically valid prediction intervals—a first for latent-space ROMs. We demonstrate this UQ pipeline on a cloud microphysics ROM trained to predict the evolution of cloud DSDs during coalescence and the formation of precipitation [dejong2025data], producing reliable UQ estimates that allow practitioners to rigorously evaluate individual components of the ROM architecture.
@@ 2 Proposed UQ framework
A latent-space dynamical ROM consists of a data space, $\mathcal{X}\subseteq\mathbb{R}^{d}$, and a latent space, $\mathcal{Z}\subseteq\mathbb{R}^{m}$, where $m\ll d$; an encoder $E:\mathcal{X}\rightarrow\mathcal{Z}$ and a decoder $D:\mathcal{Z}\rightarrow\mathcal{X}$; and a dynamical system model $F:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{Z}$ defined on the latent space, where $\mathcal{T}\subseteq\left[0,T\right]$ for some final time $T>0$. (See Figure 1.)
Figure 1: A generic latent-space dynamical model with fixed dynamical time-step $\Delta t$. Figures 1a and 1b show the reconstruction and dynamics sub-models, respectively, that comprise the end-to-end model architecture shown in Figure 1c.
We consider the setting where we observe $n$ independent time-evolving realizations of a physical system in $\mathcal{X}$: $\{x_{t}^{(j)}\}_{t\in\mathcal{T}}$ for $j=1,\dots,n$. The proposed pipeline applies regardless of whether model components are trained separately or end-to-end. Hence, at each fixed time $t\in\mathcal{T}$ and across samples $j=1,\dots,n$, we compute predictive uncertainties in the components of a latent-space dynamical model: 1) reconstruction from the latent space, 2) dynamics in the latent space, and 3) the entire pipeline combined (end-to-end), as shown in Figure 1.
@@
For DSD-valued predictions, we construct two-sided conformal prediction intervals using the $\alpha/2$ and $1-\alpha/2$ empirical quantiles of the signed residuals [barber_predictive_2021]. This approach permits asymmetric upper and lower bounds when the residual distribution is skewed—a crucial feature for DSDs, as these are non-negative and often attain values near zero, allowing prediction intervals to reflect physical uncertainty better. Meanwhile, for latent-space predictions, whose outputs are multivariate and correlated, we instead use a scalar conformity score based on the Mahalanobis distance between predicted and true latent variables, providing a covariance-aware measure of predictive error. In either case, conformal sets are computed at each timestep separately, so that uncertainty can be tracked in time alongside latent-space dynamics and DSD evolution. See Appendix B for further details.
@@ 3 Application to Cloud Microphysics ROM
We demonstrate the UQ pipeline with a cloud microphysics ROM, trained on DSDs from particle-based simulations of warm-rain coalescence, and developed in detail in [dejong2025data]. Coalescence is the process of small cloud droplets colliding and merging to form larger ones, eventually leading to large rain drops that precipitate [khain_representation_2015]. As shown in Figure 2, droplet mass shifts from smaller to larger size bins under coalescence, and the strong nonlinearity of the process drives the emergence and disappearance of multiple modes.
@@
New parameterizations of cloud microphysics must not only outperform traditional schemes in accuracy and efficiency, but also quantify structural errors and parametric uncertainties that currently hinder the accuracy of large-scale models. Nonlinear ROMs offer a more flexible alternative to bulk parameterizations, which impose restrictive modeling assumptions [kessler_distribution_1969, khairoutdinov_new_2000, seifert_double-moment_2001], and to linear latent-space ROMs, which could require inefficiently high-dimensional representations to capture DSD coalescence dynamics accurately [lin_enhanced_2025, romor_non-linear_2023, wan_evolve_2023]. Hence, to model droplet coalescence, we combine an autoencoder (AE) for nonlinear dimensionality reduction with parsimonious latent-space ODEs, based on the Sparse Identification of Nonlinear Dynamics (SINDy) technique [brunton_discovering_2016, champion_data-driven_2019, dejong2025data].
For details on the AE–SINDy architecture, training, and accuracy, see Appendix C, the associated code repository, and the companion paper by De Jong et al. [dejong2025data], which develops, trains, and evaluates the AE-SINDy surrogate in depth. The companion paper also applies the UQ framework introduced here to two additional latent-space ROMs—each using a similar autoencoder architecture but different dynamical models in the latent space—thereby highlighting broader context and applicability. Here, we focus specifically on the UQ pipeline and its expanded application to the AE–SINDy surrogate.
3.1 Dataset from LES with superdroplet method
We represent each DSD as a mass-density function $\frac{dm}{d\ln r}$, where $r$ is the droplet radius, so that the DSD is defined per logarithmic radius interval. To train and evaluate the AE-SINDy surrogate, we use PSD trajectories generated from large eddy simulations (LES) of warm-rain coalescence with a high-fidelity Lagrangian particle representation of cloud droplets [dejong2025data, shima_super-droplet_2009, unterstrasser_collisional_2020]. DSDs are averaged over a (200200m)3 cubic domain, filtered for the presence of cloud condensate, and discretized into $N_{\text{bins}}=64$ bins uniformly spaced in $\ln r$. For each simulation grid cell, a $600$s coalescence-only forward integration produces the evolution of its PSD from $t=0$ to $t=600$s at $\Delta t=10$s intervals. These trajectories provide both the instantaneous DSDs and the associated time derivatives used for training the AE-SINDy model and for assessing predictive uncertainty across different initial conditions.
3.2 Computational efficiency and UQ
The LES with Lagrangian microphysics used to generate the dataset are computationally intensive, using $3\times 10^{6}$ grid cells, each with 128 particles for a total of $\sim 10^{8}$ Lagrangian particles per simulation. Collisional-coalescence is computed via linear stochastic pairwise sampling within each grid cell [shima_super-droplet_2009], which scales linearly with the number of particles. Combined with Lagrangian advection and Eulerian-Lagrangian coupling, these simulations tend to require hundreds of CPU-hours for a single multi-hour LES. Traditional Eulerian binned microphysics can be far less expensive, but still evolves 3030100100 prognostic DSD bins per cell with quadratically scaling computations for collisional coalescence [khain_representation_2015].
In contrast, the AE-SINDy surrogate used here compresses each $64$-bin PSD into a $4$-dimensional latent state governed by an ODE system whose evaluation is $\mathcal{O}(1)$ per grid cell per timestep. This yields reductions of several orders of magnitude in computational cost relative to SDM and at least an order of magnitude reduction relative to bin microphysics.
These reductions motivate the need for rigorous UQ. The UQ pipeline introduced here identifies when the surrogate faithfully reproduces the high-fidelity Lagrangian microphysics and when surrogate uncertainty becomes dominant. Furthermore, this pipeline allows us to measure how both structural uncertainty in the AE-based compression of DSDs and parametric uncertainty in the SINDy-identified latent dynamics propagate through DSD coalescence predictions. The examples in Figure 2 illustrate prediction intervals for three representative DSDs at fixed times, while Figure 3 summarizes the uncertainty propagated across different components of the ROM as a function of time and nominal miscoverage level $\alpha$.
In Figure 3, the $y$-values report an average prediction interval width for each indicated ROM component. For DSD-valued predictions, this is computed as a normalized, total mass-weighted integral of prediction interval widths across bins, while for latent-space predictions, it is computed via the Mahalanobis distance-derived prediction ellipsoids in latent space, whose geometry is normalized by the residual covariance. Our analysis highlights the ability of the presented UQ pipeline to characterize data-driven ROMs by identifying the specific scales and processes where model improvements are most necessary.
@@ 3.2 Computational efficiency and UQ
Figure 2: Initial and final states for three sample DSD trajectories from the dataset, predicted using the trained AE-SINDy architecture (“Model”) with empirical prediction intervals provided via CV+ conformal predictions (using $k=20$ folds) at varying nominal coverage levels. These predictions are also compared with the actual DSD final states from the dataset (“Data”).
@@
Figure 3: Average prediction interval width vs. time, computed by applying vanilla CP, split CP (using a 6060-2020-2020 train-validation-test split), and CV+ conformal predictions (using $k=20$ folds), respectively, to the indicated components of the AE-SINDy pipeline at varying nominal coverage levels $1-\alpha$. Interval widths are computed using the mass-normalized integral of the area between prediction bands across bins for DSD-valued predictions and the volume of the Mahalanobis distance-derived prediction ellipsoids for latent-space predictions.
@@ 4 Results & Discussion
Figure 3 illustrates not only predictive accuracy on unseen data but also qualitative patterns of uncertainty propagation across different components of the AE-SINDy architecture. As anticipated, reconstruction uncertainty intervals, which characterize the autoencoder alone, are consistent across time. The latent dynamics exhibit rapid incipient growth in uncertainty that slows over time, reflecting the challenging cloud-to-rain transition before droplets settle into a rain-dominant coalesced state. By contrast, the full ROM produces nearly linear growth in the end-to-end predictive errors. This highlights a key advantage of component-wise uncertainty analysis in ROMs: we detected how latent errors are effectively “smoothed out” by the autoencoder, yielding linear error growth in the final predictions.
@@
While all conformal prediction methods achieved near-nominal coverage (see Appendix A), CV+ yielded wider average intervals to produce more reliable coverage, but at a higher computational cost. Although this can be done in parallel, CV+ requires retraining the surrogate model $k$ times, compared to just once for vanilla or split conformal. The cost of training the surrogate model will therefore determine practical choices for CP techniques in future applications.
@@ 5 Limitations & Future Directions
The agreement across CP methods and variability in empirical coverages was notably better for the reconstruction and end-to-end network predictions than for latent dynamical predictions (cf. Figure 3). This is likely due to greater variance in prediction fidelity across times and variables—as well as a wider range of magnitudes overall—in latent predictions compared to normalized DSD predictions. Increasing training data—for instance, by altering the initial conditions or dynamical driver of the cloud LES (see Appendix C.1 or [dejong2025data])—could reduce this variability, yielding more consistent CP intervals and improving coverage accuracy on the test set.
@@
Taken together with the companion study De Jong et al. [dejong2025data], which develops the AE-SINDy surrogate itself, this work lays the foundation for a unified framework for both rigorous UQ of latent-space ROMs and its relevance in the efficient modeling of warm-rain microphysics.
@@
This work was performed in part under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory (LLNL) under Contract DE-AC52-07NA27344 and supported by the Laboratory Directed Research and Development Program (LDRD), project number 25-ERD-045. J. Katona was partially supported by a stipend and teaching fellowship from the Yale Graduate School of Arts and Sciences. The authors have declared that none of them have any competing interests. Released under IM number LLNL-CONF-2010541.
The authors would like to thank LLNL—particularly Livermore Computing—for their support and for providing high-performance computing resources that enabled the simulations, data analyses, and figure generation in this study. The authors also appreciate the support and valuable advice of Peter Caldwell, Hassan Beydoun, Aaron Donahue, Chris Golaz, Youngsoo Choi, and many others in AEED and CASC at LLNL.
@@ Appendix A Validation of empirical prediction intervals
Tables 1 and 2 give summary statistics—means and standard deviations, and medians, respectively—for the nominal coverages in the outputs for different subsets of the network: autoencoder/reconstruction alone, latent dynamical model alone, and the entire end-to-end network predictions. These statistics are averaged over all output variables—64 bins for DSD data or 4 latent variables—and times—61 timesteps at 10-second intervals.
@@
Table 1: Empirical coverage (in %, given as empirical mean $\pm$ empirical standard deviation across all times and output coordinates) for prediction intervals at marginal coverage levels 90%, 95%, 98%, and 99%. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Empirical medians appear in Table 2.
Sub-model CP Method $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Reconstruction Vanilla 88.56 $\pm$ 3.16 93.86 $\pm$ 2.30 96.96 $\pm$ 1.65 98.00 $\pm$ 1.38
Reconstruction Split 87.70 $\pm$ 4.31 92.87 $\pm$ 3.62 96.10 $\pm$ 2.57 97.34 $\pm$ 2.17
Reconstruction CV+ 89.04 $\pm$ 3.09 94.36 $\pm$ 2.21 97.34 $\pm$ 1.60 98.28 $\pm$ 1.35
Latent dynamics Vanilla 89.38 $\pm$ 5.41 95.16 $\pm$ 3.36 97.88 $\pm$ 1.64 98.98 $\pm$ 1.04
Latent dynamics Split 88.22 $\pm$ 4.33 94.30 $\pm$ 1.87 97.32 $\pm$ 1.11 98.44 $\pm$ 1.26
Latent dynamics CV+ 95.23 $\pm$ 3.45 98.36 $\pm$ 1.93 99.44 $\pm$ 1.19 99.83 $\pm$ 0.53
End-to-end Vanilla 88.65 $\pm$ 3.35 93.73 $\pm$ 2.55 96.79 $\pm$ 1.96 97.94 $\pm$ 1.54
End-to-end Split 87.36 $\pm$ 4.57 92.79 $\pm$ 3.55 96.25 $\pm$ 2.54 97.45 $\pm$ 1.96
End-to-end CV+ 90.56 $\pm$ 3.57 95.20 $\pm$ 2.43 97.64 $\pm$ 1.68 98.46 $\pm$ 1.37
@@
Table 2: Empirical coverage (in %, given as the empirical median across all times and output coordinates) for prediction intervals at marginal coverage levels $1-\alpha=$90%, 95%, 98%, and 99%. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Corresponding means and standard deviations appear in Table 1.
Sub-model CP Method $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Reconstruction Vanilla 88.71 93.55 96.77 98.39
Reconstruction Split 87.10 93.55 95.97 97.58
Reconstruction CV+ 89.52 94.35 97.58 98.39
Latent dynamics Vanilla 91.13 95.97 97.58 99.19
Latent dynamics Split 88.71 94.35 97.58 98.39
Latent dynamics CV+ 95.97 99.19 100.00 100.00
End-to-end Vanilla 88.71 92.74 96.77 98.39
End-to-end Split 87.10 92.74 96.77 97.58
End-to-end CV+ 91.13 95.16 97.58 99.19
@@
Across all CP methods, the empirical coverage was generally close to nominal levels, indicating good calibration. While split conformal is theoretically more accurate than vanilla conformal [angelopoulos2021gentle, lei2018distribution], for both reconstruction and end-to-end outputs, vanilla and split conformal performed similarly, with mean coverages typically within 1%–2% of the target and relatively low variability and standard deviations typically under $\sim$4% (aside from when $1-\alpha=90\%$). By contrast, CV+ consistently achieved slightly higher accuracy, producing empirical coverages that were closer to the nominal rates in both mean and median, especially at higher confidence levels. This improvement was most apparent in the end-to-end model, where CV+ coverage levels tracked the nominal ones more tightly than vanilla or split conformal did.
The latent dynamical model displays somewhat different coverage behavior from what was observed for the reconstruction and end-to-end components. For vanilla and split CP, latent-space coverages are still near the nominal levels—typically within 1-2%—but with slightly larger variability across samples than for reconstructions and end-to-end outputs. In contrast, CV+ produces systematically conservative intervals for the latent dynamics, with mean coverages around 95% vs. 90% at the 90% nominal level and mean coverages exceeding 98% at the 95–99% nominal levels; in fact, the corresponding medians saturated at 100% for the two highest nominal coverages. These results suggest that CP methods measure reconstruction and end-to-end prediction uncertainties reasonably well, while for latent dynamics, the coverage tends to be more conservative, especially for CV+, which provides the most stable yet also most pessimistic intervals.
@@
To test more specifically for consistency in the predictive intervals across different CP methods, we also refer the reader to Figure 3. For all three subsets of the network, the predictive errors become less consistent as $\alpha\rightarrow 0$. This is because the convergence of empirical quantiles to their true values depends strongly on the miscoverage rate $\alpha$ [serfling1980chapter2, vaart1998chapter21]. In particular, for CP, predictive intervals at smaller $\alpha$ require larger calibration sets to stabilize because extreme quantiles converge more slowly, whereas more central quantiles yield more stable intervals with fewer samples [lei2018distribution, vovk_algorithmic_2022].
Appendix B Nonconformity scores